OpenStudy (anonymous):

26 is what percent of 40?

OpenStudy (anonymous):

26/40 = x/100 cross multiply 40x=2600 x= 65%

OpenStudy (anonymous):

is it multiple choice?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok cool is that one of the answers?

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

ok my work here is done! =)

OpenStudy (anonymous):

;-)